OpenStudy (anonymous):

Can someone please help me? Find the fifth roots of 243(cos 300° + i sin 300°).

OpenStudy (anonymous):

yeah it is much easier than you think

OpenStudy (anonymous):

first of all \(243=3^5\) so \(\sqrt[5]{243}=3\) , we will use the 3 in each answer now all you need to do is divide \(300\div 5=60\) so one answer is \[3\left(\cos(60)+i\sin(60)\right)\]

OpenStudy (anonymous):

guess we lost @Module9

OpenStudy (anonymous):

Sorry, I started working on it and just took off. I think I got it now. Thanks so much for your help.

OpenStudy (anonymous):

you still have more work to do, you have to go around the circle again and then divide by 5 again

OpenStudy (anonymous):

or you can divide the circle up in to five equal parts with one of them at \((\cos(60),\sin(60))\)

OpenStudy (anonymous):

I know, I have to do it for each root. my original issue was that the first answer was coming out all wrong, and I realized I had the theorem written down wrong.

OpenStudy (anonymous):

oh ok, good luck

OpenStudy (anonymous):

@satellite73 can you keep explaining this?